# Computation of attractor dimension and maximal sums of Lyapunov exponents using polynomial optimization

### v1

@@ 3.2 A three-degree-of-freedom Hamiltonian system
For an example where LE sums for different $k$ are maximized on different orbits, we consider a higher-dimensional system whose large number of symmetries keeps the computational cost tractable. In particular, we consider a Hamiltonian system with Hamiltonian function
@@
the first three components of $x$ being the position variables, and the final three the corresponding momenta. This is a specific form of a triaxially symmetric potential studied by Palacián et al. [35], which has applications in galactic dynamics [6, 4]. The Hamiltonian 118 gives rise to a system of ODEs in 6 variables,
$\frac{\mathrm{d}x}{\mathrm{d}t}=\begin{bmatrix}x_{4}\\ x_{5}\\ x_{6}\\ -x_{1}-2x_{1}x_{2}^{2}-2x_{1}x_{3}^{2}\\ -x_{2}-2x_{2}x_{1}^{2}-2x_{2}x_{3}^{2}\\ -x_{3}-2x_{3}x_{1}^{2}-2x_{3}x_{2}^{2}\end{bmatrix}$. (119)
The right-hand side of 119 is equivariant under $x\mapsto\Lambda x$ where $\Lambda$ is any matrix in the subgroup of $O(6)$ generated by
@@
We consider the fixed energy level $H=1$. At this value of $H$, there are no equilibria, so $\mu_{3}=\mu_{4}=0$ on all orbits. We consider only orbits whose position variables remain in the region where $\frac{1}{2}\left(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)+\left(x_{1}^{2}x_{2}^{2}+x_{2}^{2}x_{3}^{2}+x_{3}^{2}x_{1}^{2}\right)\leq 1$, depicted in figure 2, and whose full state vector $x$ remains in the compact sphere $\frac{1}{2}|x|^{2}\leq 1$. Two families of particularly short periodic orbits, which are also depicted in figure 2, were found by a brute-force search over random initial conditions. The first family, which we call $S_{1}$, has a period of approximately 5.278 and passes near
$x=(0.83871,0,0,0,0,1.13867),$
or near points that map to this $x$ under a symmetry. The $S_{1}$ orbits have LEs of approximately
@@
The second family of orbits, which we call $S_{2}$, has a period of approximately 4.857 and passes near
x=(0.57566,0.57566,0,0.39004,0.39004,0.90186),x=(-0.57566,0.57566,0,0.39004,0.39004,0.90186),
or near a symmetry-related point. These $S_{2}$ orbits have LEs of approximately
@@
Figure 2: Plots of the position variables $(x_{1},x_{2},x_{3})$ for the Hamiltonian system 118 at the energy level $H=1$. Left: The family of periodic orbits $S_{1}$. Right: The family of periodic orbits $S_{2}$. Trajectories are confined to stay within the region $\frac{1}{2}\left(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}\right)+\left(x_{1}^{2}x_{2}^{2}+x_{2}^{2}x_{3}^{2}+x_{3}^{2}x_{1}^{2}\right)\leq 1$, which is bounded by the shaded surface. On $S_{1}$, the LE sums are $M_{1}=M_{2}=M_{3}=M_{4}=M_{5}=0.42913$ and on $S_{2}$, $M_{1}=M_{5}=0.25919$ and $M_{2}=M_{3}=M_{4}=0.51837$.
$d_{x}$ $d_{z}$ $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $k=6$
2 0 1.2361 2.0232 2.1082 2.0232 1.2361 0.0000
2 2 0.4843 0.9216 0.9216 0.4843 0.0000
4 0 1.1855 2.0046 2.0611 2.0046 1.1855 0.0000
4 2 0.4292 0.4292 0.0000
@@
Table 3: Upper bounds on the sum $M_{k}$ of the $k$ leading LEs, computed by solving the SOS optimization problem in proposition 2.15 for the Hamiltonian system 119. Polynomial degrees are at most $d_{x}$ in $x$ and $d_{z}$ in $z^{k}$. Tabulated values are rounded up to the precision shown, and the underlines indicate the digits that agree (after rounding) with the values computed by numerical integration for the periodic orbits $S_{1}$ ($k=1,5,6$) or $S_{2}$ ($k=2,3,4$) of figure 2. Values are missing in cases where computations were prohibitively memory-intensive.
$d_{x}$ $k=1$ $k=2$ $k=3$ $k=4$ $k=5$ $k=6$
2 0.4311 0.6141 0.7579 0.6141 0.4311 0.0000
4 0.4292 0.5184 0.5711 0.5184 0.4292 0.0000
6 0.4292 0.5184 0.5300 0.5184 0.4292 0.0000
Table 4: Upper bounds on the sum $M_{k}$ of the $k$ leading LEs, computed by solving the SOS optimization problem in proposition 2.17 for the Hamiltonian system 119. Polynomial degrees are at most $d_{x}$ in $x$ and 2 in $w^{k}$. The values in this table are not rounded. The underlines indicate the digits that agree (after rounding) with the values computed by numerical integration for the periodic orbits $S_{1}$ ($k=1,5,6$) or $S_{2}$ ($k=2,3,4$) of figure 2.
Because this system is Hamiltonian, volumes in state space are conserved, and our methods for bounding Lyapunov dimension are not relevant. We carried out SOS computations to bound LE sums using the sphere projection and shifted spectrum approaches, and tables 3 and 4 give the results. In both cases we add the explicit constraint $|x|^{2}\leq 2$ along with $H(x)=1$ in the definition of the set $\mathcal{B}$ because this improves the numerical results. Using the shifted spectrum approach, we find sharp bounds for $M_{1}$, $M_{2}$, $M_{4}$, $M_{5}$ and $M_{6}$. Using the sphere projection approach, we found sharp bounds on $M_{1}$, $M_{5}$ and $M_{6}$. Sharp bounds for the other $k$ values seemed to require larger polynomial degrees that were beyond our computational resources.
Given the Hamiltonian structure of the system, bounding $M_{5}$ is directly equivalent to bounding $M_{1}$, bounding $M_{4}$ is equivalent to bounding $M_{2}$, and bounding $M_{6}$ is trivial, since this is always zero. These facts are borne out by the results. Bounding the sum $M_{3}$ is not a priori identical to any other case, but the lack of equilibria means that we expect the true value to be equal to $M_{2}$ and $M_{4}$. Furthermore, bounds for $M_{3}$ require a particularly high-dimensional augmented system of dimension $n+n_{k}=26$, which is computationally challenging at even modest polynomial degree despite exploiting all symmetries of the system. This is especially true for the sphere projection method, which has higher polynomial degree in the tangent space dynamics. In table 3, the missing values were not possible to compute when solving SDPs with Mosek due to insufficient memory, despite having 3TB of memory. This could be surmounted by using a first-order SDP solver at the expense of very long runtimes and/or lower precision. In table 4, the $M_{3}$ result is not sharp at the maximum degree $d_{x}=6$ that was computationally feasible.
@@ 4 Conclusion
We have described two distinct approaches to bounding sums of Lyapunov exponents as well as Lyapunov dimension in ODE systems. The general versions of our formulations require optimizing over auxiliary functions in the class $C^{1}$. We prove that most of these formulations give sharp bounds under weak assumptions, but the optimization problems are not tractable in general. In the case of ODEs with polynomial right-hand sides, our bounding formulations can be relaxed using SOS constraints, which we again prove are convergent, and which are often computationally tractable using SOS optimization. Carrying out such SOS computations for particular ODEs, we obtained nearly perfect bounds in challenging examples. For one of the examples, the sphere projection method gave much better conditioned SDPs in practice. For the other example, the shifted spectrum method was more appropriate due to onerous memory requirements of the sphere projection method. In section 3.1, we demonstrated how the results of our bounding methods could be used to find the optimal, potentially previously undetected, periodic orbits which maximize the bounds.
All of our methods could be used to construct fully rigorous bounds via a computer assisted proof, making use of interval arithmetic or rational projection [12, 37]. In both of our examples, the results of the SOS programs allowed us to find periodic orbits that attained our bounds to within several digits, thus confirming their sharpness. In principle, computer-assisted proofs based on our framework could also give rigorous upper bounds on LEs in partial differential equations, by combining our methods with some in Goluskin and Fantuzzi [13], but such computations would be challenging.
In the thesis of Oeri [33], an analogous approach to the sphere projection method was formulated for bounding the leading LE of discrete-time dynamical systems, with an SOS implementation for polynomial maps. This can be directly adapted to sums of LEs by using additive (rather than multiplicative) compound matrices. An approach analogous to our shifted spectrum method is also possible for bounding sums of LEs for discrete-time systems, and methods can additionally be derived for the Lyapunov dimension. In general, these are computationally difficult because the equivalent of the Lie derivative is a composition of the dynamics with an auxiliary function, leading to very high polynomial degrees. This may be the focus of a future publication.
@@
Code to reproduce all the examples in this paper is available at https://github.com/jeremypparker/Parker_Goluskin_LEs.
@@ Acknowledgements
The authors thank Samuel Punshon-Smith and Anthony Quas for useful discussions. During this work DG was supported by the NSERC Discovery Grants Program (awards RGPIN-2018-04263 and RGPIN-2025-06823). Computational resources were provided by the Digital Research Alliance of Canada.
@@ Appendix A Exterior products and compound matrices
The exterior (or wedge or Grassmann) product is a multilinear map of vectors, defined by the property that $x_{2}\wedge x_{1}=-x_{1}\wedge x_{2}$, and $x_{1}\wedge x_{2}=0$ only if $x_{1}=\alpha x_{2}$ for some $\alpha$, or $x_{2}=0$. An exterior product of $k$ vectors in a vector space of dimension $n\geq k$ is called a $k$-vector. The properties of the exterior product imply that permuting the $x_{i}$ in $x_{1}\wedge\dots\wedge x_{k}$ either leaves the $k$-vector unchanged or negates it, and that the $k$-vector is nonzero if and only if the $x_{i}$ are linearly independent. The set of $k$-vectors in $\mathbb{R}^{n}$ is therefore a linear space of dimension $n_{k}=\tfrac{n!}{(n-k)!}$. We identify this space with $\mathbb{R}^{n_{k}}$ and choose as a basis the set of $k$-vectors of the form
$e_{i_{1}}\wedge\dots\wedge e_{i_{k}},$ (120)
where $i_{1}<\ldots<i_{k}$ and $e_{i}$ is the $i^{th}$ standard basis vector in $\mathbb{R}^{n_{k}}$. The standard Euclidean norm on this space satisfies the property that
$|x_{1}\wedge\dots\wedge x_{k}|$ (121)
is the volume of the parallepiped spanned by $x_{1},\dots,x_{k}\in\mathbb{R}^{n}$ [47].
Linear maps on vectors in $\mathbb{R}^{n}$ induce linear maps on corresponding $k$-vectors in $\mathbb{R}^{n_{k}}$. Likewise, linear ODEs on $\mathbb{R}^{n}$ induce linear ODEs on $k$-vectors. As defined below, the map on $k$-vectors induced by a map $A$ is captured by the multiplicative compound $A^{(k)}$, and the ODE for $k$-vectors induced by an ODE with right-hand side $A$ is captured by the additive compound $A^{[k]}$. The multiplicative compound is common in many different areas of mathematics, and is known variously as the compound representation, exterior power representation, wedge power representation or simply the induced action of a linear map. Compound matrices have been applied previously in the calculation and bounding of LEs [26, 29, 30, 18] and also stability of dynamical systems [31]. Here we state key definitions and theorems; a more detailed introduction can be found in Marshall et al. [28, Chapter 19F], in more generality for complex non-square matrices.
@@
    The $k^{th}$ multiplicative compound of $A$ is the real matrix $A^{(k)}\in\mathbb{R}^{{n_{k}}\times n_{k}}$ whose entries are the $k\times k$ subdeterminants of $A$, taken in lexicographic order.
@@
    The $k^{th}$ additive compound of $A$ is the real matrix $A^{[k]}\in\mathbb{R}^{{n_{k}}\times n_{k}}$ such that
@@
The lexicographic order in A.1 refers to the ordering of all $n_{k}$ permutations of $k$ increasing integers chosen from $\{1,\ldots,n\}$. As an example, the lexicographic order of $k=3$ integers in the case $n=4$ is
@@
where $a_{p,q}$ denotes the $(p,q)$ element of $A$. Note that the row and column indices come from the second and third permutations, respectively. It follows from the definitions that $A^{(1)}=A^{[1]}=A$, and $A^{(n)}=\det{A}$ whereas $A^{[n]}=\mathrm{tr}A$. In fact, the eigenvalues of multiplicative and additive compounds are respectively the products and sums of the eigenvalues of the original matrix, which gives rise to their names. An explicit example of additive compound matrices is given in 111.
The next two propositions list additional properties that will be useful to us for multiplicative and additive compounds, respectively. In proposition A.2, part 1 is a special case of the Cauchy–Binet theorem. Parts 2–4 follow quickly from the definition, and the last part is theorem 6.2.10 of Fiedler [10].
@@
    For $k$-vectors identified with $\mathbb{R}^{n_{k}}$ and for $A\in\mathbb{R}^{n\times n}$,
@@
    $A^{[k]}\left(x_{1}\wedge\dots\wedge x_{k}\right)=(Ax_{1}\wedge x_{2}\wedge\dots\wedge x_{k})+(x_{1}\wedge Ax_{2}\wedge\dots\wedge x_{k})+\dots+(x_{1}\wedge x_{2}\wedge\dots\wedge Ax_{k}).$
@@
    If $Q$ is invertible, $\left(QAQ^{-1}\right)^{[k]}=Q^{(k)}A^{[k]}\left(Q^{(k)}\right)^{-1}$.

### v2

@@ 3.2 A three-degree-of-freedom Hamiltonian system
For an example in which LE sums $M_{k}$ for different $k$ are not all maximized on the same orbits, we consider a 6-dimensional ODE whose numerous symmetries keep the computational costs tractable. In particular, we consider a Hamiltonian system whose Hamiltonian function is
@@
where the first three components of $x\in\mathbb{R}^{6}$ are position variables, and the other three components are corresponding momenta. This is a specific form of a triaxially symmetric potential studied by Palacián et al. [36] that has applications in galactic dynamics [7, 3]. The Hamiltonian 114 gives rise to the ODE system
$\frac{\mathrm{d}}{\mathrm{d}t}x=f(x),\quad\text{where}\quad f(x)=\begin{bmatrix}x_{4}\\ x_{5}\\ x_{6}\\ -x_{1}-2x_{1}x_{2}^{2}-2x_{1}x_{3}^{2}\\ -x_{2}-2x_{2}x_{1}^{2}-2x_{2}x_{3}^{2}\\ -x_{3}-2x_{3}x_{1}^{2}-2x_{3}x_{2}^{2}\end{bmatrix}$. (115)
We consider dynamics only on the $H=1$ energy surface, where there are no equilibria. Our semialgebraic specification of this domain also includes the redundant constraint that $|x|^{2}\leq 2$, so it is
$\mathcal{B}=\bigg\{x\in\mathbb{R}^{3}\;:\;g(x)\geq 0,~h(x)=0\bigg\},$ (116)
where
$g(x)=2-|x|^{2},\quad h(x)=H(x)-1.$ (117)
Adding the inequality constraint does not change the set $\mathcal{B}$, but it changes the SOS Positivestelansätze that certify nonnegativity on $\mathcal{B}$, and we find empirically that this gives better bounds at lower computational cost in this example. The ODE 115 and its domain 116 are both invariant under $x\mapsto\Lambda x$ for all $\Lambda$ in the 48-element octahedral symmetry group. Represented as a subgroup of $O(6)$, this group is generated by a sign-change $\Lambda_{1}$ and permutations $\Lambda_{2}$ and $\Lambda_{3}$, where
@@
Various relations between the LEs hold a priori because this system is Hamiltonian and has no equilibria. Hamiltonian systems conserve state space volume, meaning that $\mathrm{tr}(Df)=0$ pointwise. This implies that $M_{6}=0$ on all trajectories, and that there are no attractors, so our methods for bounding Lyapunov dimension are not relevant. Hamiltonian structure also implies that LEs come in pairs of $\mu_{k}$ with opposite sign [48], so $(\mu_{1},\mu_{2},\mu_{3})=(-\mu_{6},-\mu_{5},-\mu_{4})$. Since there are no equilibria, all trajectories have a zero LE corresponding to time translation, which in Hamiltonian systems guarantees at least two zero LEs, here meaning $\mu_{3}=\mu_{4}=0$.
Relations between the LE sums $M_{k}$ follow from the $\mu_{k}$ relations, which imply
$M_{1}=M_{5}=\mu_{1},\quad M_{2}=M_{3}=M_{4}=\mu_{1}+\mu_{2},\quad M_{6}=0.$ (118)
Therefore, to obtain sharp upper bounds on the maxima of all LE sums, it suffices to compute sharp bounds on $M_{1}$ and $M_{2}$. Nevertheless, in order to test our methods, we computed bounds for all $M_{k}$ aside from $M_{6}$. The fact that $M_{6}=0$ on all trajectories is immediate since $Df^{[6]}=0$ pointwise, and simply taking $V=0$ in our bounding formulations would give $M_{6}\leq 0$.
We have computed a number of unstable periodic orbits lying in $\mathcal{B}$. The maxima of $M_{1}$ and $M_{2}$ among these orbits provide candidates for the maxima over all trajectories in $\mathcal{B}$, but this can only be confirmed by the upper bounds on $M_{k}$ reported below. The periodic orbits were found by a brute-force search over random initial conditions, uniformly distributed over the unit box in $(x_{1},x_{2},x_{4},x_{5})$ on the $x_{3}=0$ Poincaré section, with $x_{6}$ chosen so that $H(x)=1$. These were converged by a Newton-like method to periodic orbits that close after one iteration of the Poincaré return map. Among the periodic orbits we found, $M_{1}$ is maximized on each orbit in a set of 6 symmetry-related orbits that we call $S_{1}$, while $M_{2}$ is maximized on each in a set of 4 symmetry-related orbits that we call $S_{2}$. Figure 2 shows the $S_{1}$ and $S_{2}$ orbits plotted in position space.
Figure 2: Position coordinates $(x_{1},x_{2},x_{3})$ for the Hamiltonian system 114 showing the $S_{1}$ (left) and $S_{2}$ (right) families of periodic orbits (thick lines). The $S_{1}$ orbits have LE sums $M_{1}=M_{2}=M_{3}=M_{4}=M_{5}=0.42913$, and the $S_{2}$ orbits have $M_{1}=M_{5}=0.25919$ and $M_{2}=M_{3}=M_{4}=0.51837$. The thin lines depict the part of the $H=1$ surface with zero momenta, which is the boundary of possible position coordinates.
The $S_{1}$ orbits have a period of approximately 5.278 and pass near one of the symmetry-related points $\Lambda x$ for
$x=(0.83871,0,0,0,0,1.13867).$
Their LEs are approximately
@@
The $S_{2}$ orbits have a period of approximately 4.857 and pass near $\Lambda x$ for
x=(0.57566,0.57566,0,0.39004,0.39004,0.90186).x=(-0.57566,0.57566,0,0.39004,0.39004,0.90186).
Their LEs are approximately
@@
One may then guess that the maximum of $M_{1}=M_{5}$ among all trajectories in $\mathcal{B}$ is close to 0.42913, and that the maximum of $M_{2}=M_{3}=M_{4}$ is close to 0.57566. These guesses are confirmed by the upper bounds that we report next.
As in the example of section 3.1, we formulate SOS optimization problems by choosing finite-dimensional spaces over which to optimize $V$ and the other polynomials appearing in Positivstellansätze. These spaces enforce symmetries of the problem (cf. section 2.5) and maximum polynomial degrees (cf. the fourth paragraph of section 3). We computed bounds on $M_{k}$ using the SOS programs 80 and 91 from the sphere projection and shifted spectrum approaches, respectively. In both cases the total number of variables in the SOS program for $k=1,2,3,4,5$ are $6+C_{6}^{k}=12,21,26,21,12$, respectively. Due to high computational cost with this number of variables, the degree of $V$ in the $C_{6}^{k}$ tangent space variables was no larger than 2. The SOS constraints involve the additive compounds of the Jacobian matrix of $f$ in 115. For brevity we do not write down these compounds, as we did in the example of section 3.1. They can be found in our included code, where they are constructed using the Julia package DynamicPolynomials.jl [22].
Table 2 reports bounds on $M_{k}$ computed using our two approaches. In both approaches, we vary the degree $\nu_{x}$ of $V$ in $x$, and in the sphere projection approach we also vary its degree $\nu_{z}$ in $z^{k}$.
@@
Table 2: Upper bounds on the LE sums $M_{k}$ for the Hamiltonian system 115, computed with the sphere projection and shifted spectrum approaches by solving the SOS problems in propositions 2.16 and 2.18, respectively. Polynomial degrees in $x$ are at most $\nu_{x}$. In the sphere projection approach, degrees in $z^{k}$ are at most $\nu_{z}$, and tabulated values are rounded up to the precision shown. In the shifted spectrum, expressions are quadratic in $w$, the tabulated values are the exact $B$ whose feasibility was verified, and feasibility could not be verified when the final digit of $B$ was decreased by 1. Underlines indicate digits that agree (after rounding) with the numerically computed values $M_{1}=M_{5}\approx 0.42913$ on the $S_{1}$ periodic orbits or $M_{2}=M_{3}=M_{4}\approx 0.51837$ on the $S_{2}$ orbits (cf. figure 2). Bounds are missing in cases where solution of SDPs by interior-point algorithms required more than 3TB of memory.
$\nu_{x}$ $\nu_{z}$ $M_{1}$ $M_{2}$ $M_{3}$ $M_{4}$ $M_{5}$
Upper bounds (sphere projection approach)
2 0 1.2361 2.0232 2.1082 2.0232 1.2361
2 0.4843 0.9216 0.9480 0.9216 0.4843
4 0 1.1855 2.0046 2.0611 2.0046 1.1855
2 0.4292 0.4292
Upper bounds (shifted spectrum approach)
2 0.4311 0.6141 0.7579 0.6141 0.4311
4 0.4292 0.5184 0.5711 0.5184 0.4292
6 0.4292 0.5184 0.5300 0.5184 0.4292
Using the sphere projection, bounds on $M_{1}$ and $M_{5}$ are sharp to 4 digits with $V$ of maximum degrees $(\nu_{x},\nu_{z})=(4,2)$. For $k=2,3,4$, where the SOS problems have more variables, we did not obtain sharp bounds using sphere projection. With degrees $(\nu_{x},\nu_{z})=(4,2)$ at these $k$, we could not solve the necessary SDPs using Mosek because the 3TB of memory available to us was insufficient. This could be surmounted by using a first-order algorithm, rather than an interior-point algorithm, at the expense of very long runtimes and/or lower precision. Using the shifted spectrum approach, however, we obtain bounds that are sharp to 4 digits on all $M_{k}$ aside from $k=3$. Bounds in the $k=3$ case are sharp to only 2 digits. This might be due to numerical inaccuracy in the solution of the SDP, which is larger than in any other case, or it might indicate a need for computations with $\nu_{x}=8$. Nonetheless, we remind the reader that we computed bounds with $k=3,4,5$ only to test our computational formulations. Due to the Hamiltonian structure of this example, sharp bounds on $M_{1}$ and $M_{2}$ immediately imply sharp bounds on the other $M_{k}$ also. Therefore, we have indeed obtained bounds on all $M_{k}$ that are sharp to 4 digits.
@@ 4 Conclusion
We have described two distinct approaches for computing upper bounds on sums of Lyapunov exponents and on Lyapunov dimension in ODE systems. One we call the sphere projection approach, which leads to a deterministic version of the Furstenberg–Khasminskii formula from stochastic dynamics. The other we call the shifted spectrum approach. In both approaches, our most general formulations require optimizing over auxiliary functions in the class of $C^{1}$ functions. We have proved that most of these formulations give sharp bounds under mild assumptions, but the optimization problems are not tractable in general. In the case of ODEs with polynomial right-hand sides, our bounding formulations can be relaxed using SOS constraints, yielding SOS optimization problems that are often computationally tractable, and which often converge to sharp bounds as polynomial degrees are raised. By carrying out such SOS computations for particular ODEs, we obtained nearly perfect bounds in challenging examples. For the Duffing oscillator example, the sphere projection computations had much better numerical conditioning. For the Hamiltonian example, the shifted spectrum computations were easier because the sphere projection computations had onerous memory requirements. In both examples, we confirmed sharpness of our best bounds by finding periodic orbits that saturate the bounds to at least 4 digits. We also demonstrated, in the Duffing oscillator example, how the results of our bounding computations can be used to find previously undetected orbits that saturate the bounds.
All of our computational methods could be used to construct fully rigorous bounds via computer-assisted proofs, making use of interval arithmetic or rational projection [14, 39, 40]. In principle, computer-assisted proofs based on our framework could also give rigorous upper bounds on LEs in partial differential equations, by combining our methods with some in Goluskin and Fantuzzi [13], but such computations would be challenging.
A clear future direction is to adapt our present methods to discrete-time dynamical systems. A formulation analogous to the sphere projection approach was formulated for bounding the single leading LE of discrete-time systems in the thesis of Oeri [35], including computational formulations with SOS optimization in the case of polynomial maps. This can be directly extended to sums of LEs by using multiplicative compound matrices, as opposed to the additive compounds we have used for ODE dynamics. Formulations analogous to our shifted spectrum approach are also possible for discrete-time systems. For both approaches, the SOS computations arising for polynomial maps are generally harder than the SOS computations for polynomial ODEs that we have carried out here. The reason is that the auxiliary function $V$ is composed with the map in the discrete-time case, leading to high polynomial degrees. Explanation of this and other differences in the discrete-time case is left for a future publication.
@@
Our code available at github.com/jeremypparker/Parker_Goluskin_LEs was used to carry out all computational examples.
@@ Acknowledgements
The authors thank Samuel Punshon-Smith and Anthony Quas for useful discussions, and James Meiss for suggesting possible Hamiltonian examples. During this work DG was supported by the NSERC Discovery Grants Program (awards RGPIN-2018-04263 and RGPIN-2025-06823). Computational resources were provided by the Digital Research Alliance of Canada.
@@ Appendix A Exterior products and compound matrices
The exterior algebra on $\mathbb{R}^{n}$, also called the Grassman algebra, is a graded algebra with a product that is denoted by $\wedge$ and called the exterior product. An exterior product of $k$ vectors in $\mathbb{R}^{n}$ is called a $k$-blade, and the space spanned by these is the space of $k$-vectors. The properties of the exterior product imply that permuting the $x_{i}$ in $x_{1}\wedge\dots\wedge x_{k}$ either leaves the $k$-blade unchanged or negates it, and that the $k$-blade is nonzero if and only if the $x_{i}$ are linearly independent. The set of $k$-vectors over $\mathbb{R}^{n}$ is therefore a linear space of dimension $C_{n}^{k}=\tfrac{n!}{k!(n-k)!}$. We identify this space with $\mathbb{R}^{C_{n}^{k}}$ by choosing a basis, which is the set of $k$-blades of the form $e_{i_{1}}\wedge\dots\wedge e_{i_{k}}$, where $e_{i}$ denotes the $i^{\mathrm{th}}$ standard basis vector in $\mathbb{R}^{n}$, and the indices $i_{1}<\ldots<i_{k}$ are all possible tuples of $k$ increasing integers between 1 and $n$. The standard Euclidean norm on this space satisfies the property that $|x_{1}\wedge\dots\wedge x_{k}|$ is the volume of the parallelepiped spanned by $x_{1},\dots,x_{k}\in\mathbb{R}^{n}$ [51].
Any linear map $A$ on vectors in $\mathbb{R}^{n}$ induces a linear map on corresponding $k$-vectors in $\mathbb{R}^{C_{n}^{k}}$, called the multiplicative compound $A^{(k)}$. Likewise, any linear ODE on $\mathbb{R}^{n}$ with right-hand side $A$ induces a linear ODE on $k$-vectors whose right-hand side is called the additive compound $A^{[k]}$. The multiplicative compound arises in many areas of mathematics and is known variously as the compound representation, exterior power representation, wedge power representation or simply the induced action of a linear map. Compound matrices have been applied previously to dynamical systems in the calculation and bounding of LEs [27, 29, 30, 19] and in stability analysis [32]. Here we state key definitions and theorems; a more detailed introduction can be found in Marshall et al. [28, Chapter 19F], in more generality for complex non-square matrices.
@@
    The $k^{\mathrm{th}}$ multiplicative compound of $A$ is the matrix $A^{(k)}\in\mathbb{R}^{{C_{n}^{k}}\times C_{n}^{k}}$ whose entries are the $k\times k$ subdeterminants of $A$, taken in lexicographic order.
@@
    The $k^{\mathrm{th}}$ additive compound of $A$ is the matrix $A^{[k]}\in\mathbb{R}^{{C_{n}^{k}}\times C_{n}^{k}}$ such that
@@
The lexicographic order in A.1 refers to the ordering of all $C_{n}^{k}$ permutations of $k$ increasing integers chosen from $\{1,\ldots,n\}$. In this ordering, tuples are ordered by their first element, those with the same first element are ordered by their second element, and so on. As an example, the lexicographic order of $k=3$ integers in the case $n=4$ is
@@
Note that the row and column indices come from the second and third permutations in the lexicographic ordering, respectively. It follows from the definitions that $A^{(1)}=A^{[1]}=A$, while $A^{(n)}=\det{A}$ and $A^{[n]}=\mathrm{tr}A$. The eigenvalues of the $k^{\mathrm{th}}$ multiplicative and additive compounds are the products and sums, respectively, of the $k$ eigenvalues of $A$, which is why the compounds are called ‘multiplicative’ and ‘additive’. Additive compounds for all $k$ in an explicit example with $n=4$ are given in 106.
The next two propositions give properties of multiplicative and additive compounds, respectively, that will be useful to us. In proposition A.2, part 1 is a special case of the Cauchy–Binet theorem, parts 2–4 follow quickly from the definition, and part 5 is theorem 6.2.10 of Fiedler [11]. For $A^{(k)}$ and $A^{[k]}$ to be linear mappings on $k$-blades, as in part 5 of proposition A.2 and part 2 of proposition A.3, we identify the space of $k$-vectors with $\mathbb{R}^{C_{n}^{k}}$.
@@
    For any $k$-blade $x_{1}\wedge\dots\wedge x_{k}$,
@@
    For any $k$-blade $x_{1}\wedge\dots\wedge x_{k}$,
    $A^{[k]}\left(x_{1}\wedge\dots\wedge x_{k}\right)=(Ax_{1}\wedge x_{2}\wedge\dots\wedge x_{k})+(x_{1}\wedge Ax_{2}\wedge\dots\wedge x_{k})+\dots+(x_{1}\wedge x_{2}\wedge\dots\wedge Ax_{k}).$ (123)
@@
    If $B$ is invertible, $\left(BAB^{-1}\right)^{[k]}=B^{(k)}A^{[k]}\left(B^{(k)}\right)^{-1}$.